# Multiple Choice Learning of Low Rank Adapters for Language Modeling

### v1

@@ E.2.1 Experimental setup
Architecture and training. We used the Instructed version of Qwen-2-Audio [6] as the base model, which features $\sim 8.4$ billion parameters. We trained using bfloat16 precision. We used LoRA adapters applied to the $Q,K,V$ linear projections of the attention modules, and the upside and downside projections of the feedforward blocks, for all the transformer blocks, which include both the audio encoder and language model decoder. We used a rank $r$, with $r=\alpha=8$ unless otherwise stated, with dropout equal to $0.1$. We trained with a batch size of $2$, with AdamW optimizer [40] (with $\beta_{1}=0.9$, and $\beta_{2}=0.98$), weight decay of $0.05$, using a cosine scheduler with minimum learning rate of $10^{-6}$ and maximum learning rate of $10^{-5}$, with a warmup ratio of $0.1$. Gradient clipping is used with a maximum gradient norm of $1.0$. The validation loss was computed once every epoch.
E.2.2 Additional results
Figure 5: Quality vs. Diversity on Audio Captioning with $4$ candidates.
Figure 6: Quality vs. Diversity on Audio Captioning with $3$ candidates.
Varying the number $K$ of hypotheses. Figures 5 and 6 present the same scatter plots as Figure 2, but with 4 and 3 candidates respectively. The observed trends remain consistent: LoRA-MCL continues to demonstrate an excellent balance between quality and diversity. Notably, the case with $\varepsilon=0$ performs particularly well, which we attribute to improved stability in vanilla WTA when the number of heads is reduced. Note that the detailed results with $K=5$ associated with Figure 2 are given in Tables 5 and 6 for Clotho and AudioCaps, respectively.
Table 5: Results on Clotho-V2 with 5 candidates and MAP Decoding. ‘BS’ and ‘DBS’ stand for beam search and diverse beam search respectively.
@@ E.2.2 Additional results
Training Decoding Beam Div1 Div2 mBLEU-4 BLEU-1 BLEU-4 METEOR ROUGE-L sBERT CIDEr-D SPICE SPIDEr
LoRA-MLE ($r=8$) BS 5 0.298 0.360 0.827 0.651 0.133 0.229 0.443 0.575 0.621 0.175 0.391
LoRA-MLE ($r=40$) BS 5 0.300 0.365 0.819 0.619 0.105 0.213 0.422 0.559 0.556 0.166 0.354
LoRA-MLE ($r=8$) BS 10 0.310 0.386 0.787 0.650 0.139 0.233 0.450 0.576 0.634 0.181 0.400
LoRA-MLE ($r=40$) BS 10 0.314 0.393 0.780 0.623 0.113 0.219 0.426 0.561 0.564 0.170 0.360
LoRA-MLE ($r=8$) BS 25 0.316 0.402 0.764 0.653 0.143 0.236 0.453 0.580 0.639 0.184 0.404
LoRA-MLE ($r=40$) BS 25 0.324 0.411 0.750 0.629 0.121 0.222 0.433 0.563 0.580 0.171 0.368
LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 5 0.421 0.560 0.524 0.679 0.145 0.241 0.467 0.597 0.674 0.193 0.421
LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 5 0.422 0.558 0.524 0.652 0.122 0.224 0.445 0.582 0.601 0.176 0.378
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 5 0.451 0.606 0.446 0.683 0.149 0.242 0.470 0.600 0.678 0.194 0.423
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 5 0.454 0.607 0.444 0.659 0.125 0.227 0.450 0.585 0.612 0.179 0.384
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 5 0.466 0.627 0.413 0.685 0.147 0.243 0.471 0.602 0.683 0.195 0.426
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 5 0.470 0.628 0.414 0.660 0.124 0.228 0.451 0.588 0.614 0.180 0.386
LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 10 0.376 0.475 0.722 0.667 0.139 0.238 0.459 0.593 0.656 0.190 0.413
LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 10 0.377 0.480 0.716 0.642 0.115 0.223 0.439 0.575 0.589 0.179 0.374
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 10 0.386 0.491 0.708 0.668 0.139 0.238 0.459 0.595 0.652 0.191 0.412
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 10 0.390 0.495 0.704 0.640 0.115 0.223 0.437 0.577 0.585 0.179 0.371
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 10 0.392 0.495 0.704 0.664 0.138 0.240 0.455 0.596 0.655 0.192 0.413
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 10 0.395 0.498 0.704 0.641 0.112 0.224 0.437 0.578 0.583 0.182 0.372
LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 25 0.304 0.369 0.813 0.650 0.134 0.229 0.445 0.574 0.625 0.176 0.393
LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 25 0.305 0.370 0.812 0.619 0.106 0.214 0.423 0.558 0.555 0.166 0.354
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 25 0.302 0.366 0.818 0.649 0.133 0.228 0.443 0.574 0.620 0.176 0.391
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 25 0.304 0.369 0.814 0.619 0.106 0.214 0.423 0.558 0.556 0.166 0.354
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 25 0.302 0.365 0.819 0.648 0.133 0.229 0.444 0.574 0.620 0.177 0.391
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 25 0.305 0.369 0.814 0.619 0.105 0.214 0.423 0.558 0.556 0.167 0.355
LoRA-MCL ($\varepsilon=0.0$) BS 1 0.490 0.633 0.389 0.662 0.127 0.231 0.453 0.597 0.643 0.186 0.401
LoRA-MCL ($\varepsilon=0.0005$) BS 1 0.464 0.609 0.433 0.672 0.132 0.233 0.459 0.597 0.649 0.185 0.404
LoRA-MCL ($\varepsilon=0.05$) BS 1 0.471 0.623 0.405 0.677 0.141 0.238 0.463 0.600 0.672 0.191 0.419
LoRA-MCL ($\varepsilon=0.0$) BS 2 0.477 0.622 0.407 0.675 0.147 0.237 0.468 0.598 0.670 0.191 0.417
LoRA-MCL ($\varepsilon=0.0005$) BS 2 0.449 0.594 0.453 0.687 0.151 0.243 0.471 0.597 0.679 0.193 0.424
LoRA-MCL ($\varepsilon=0.05$) BS 2 0.448 0.598 0.450 0.691 0.157 0.246 0.478 0.600 0.700 0.197 0.436
LoRA-MCL ($\varepsilon=0.0$) BS 5 0.473 0.618 0.420 0.680 0.153 0.242 0.471 0.599 0.689 0.194 0.429
LoRA-MCL ($\varepsilon=0.0005$) BS 5 0.443 0.585 0.474 0.688 0.154 0.246 0.475 0.598 0.684 0.197 0.429
LoRA-MCL ($\varepsilon=0.05$) BS 5 0.445 0.592 0.465 0.691 0.165 0.249 0.482 0.601 0.708 0.199 0.442
@@
Table 6: Results for AudioCaps with 5 candidates and MAP Decoding.
Training Decoding Beam Div1 Div2 mBLEU-4 BLEU-1 BLEU-4 METEOR ROUGE-L sBERT CIDEr-D SPICE SPIDEr
LoRA-MLE ($r=8$) BS 5 0.238 0.303 0.881 0.601 0.145 0.304 0.490 0.682 0.582 0.239 0.400
LoRA-MLE ($r=40$) BS 5 0.232 0.298 0.888 0.576 0.134 0.296 0.472 0.679 0.510 0.234 0.362
LoRA-MLE ($r=8$) BS 10 0.252 0.331 0.849 0.606 0.156 0.309 0.497 0.686 0.608 0.242 0.414
LoRA-MLE ($r=40$) BS 10 0.246 0.324 0.861 0.581 0.147 0.300 0.479 0.681 0.520 0.238 0.369
LoRA-MLE ($r=8$) BS 25 0.263 0.352 0.825 0.604 0.163 0.310 0.500 0.687 0.616 0.245 0.418
LoRA-MLE ($r=40$) BS 25 0.258 0.350 0.830 0.583 0.156 0.304 0.483 0.685 0.531 0.240 0.373
LoRA-MLE ($r=8$) DBS ($\lambda=0.2$) 5 0.312 0.454 0.684 0.620 0.166 0.317 0.508 0.697 0.622 0.257 0.421
LoRA-MLE ($r=40$) DBS ($\lambda=0.2$) 5 0.310 0.462 0.676 0.595 0.157 0.314 0.492 0.697 0.534 0.255 0.376
LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 5 0.358 0.545 0.531 0.636 0.188 0.327 0.521 0.707 0.664 0.270 0.445
LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 5 0.354 0.551 0.528 0.610 0.183 0.324 0.506 0.707 0.577 0.267 0.400
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 5 0.377 0.576 0.483 0.640 0.193 0.330 0.521 0.708 0.672 0.270 0.447
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 5 0.374 0.585 0.473 0.614 0.185 0.324 0.509 0.708 0.588 0.269 0.405
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 5 0.386 0.590 0.460 0.638 0.190 0.329 0.521 0.709 0.668 0.269 0.444
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 5 0.382 0.598 0.452 0.612 0.181 0.323 0.505 0.709 0.581 0.266 0.401
LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 10 0.314 0.439 0.777 0.624 0.180 0.318 0.510 0.699 0.636 0.262 0.431
LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 10 0.309 0.440 0.780 0.605 0.170 0.315 0.496 0.699 0.558 0.260 0.390
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 10 0.320 0.448 0.770 0.624 0.176 0.318 0.507 0.700 0.626 0.263 0.426
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 10 0.316 0.451 0.771 0.604 0.167 0.313 0.496 0.698 0.556 0.258 0.388
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 10 0.325 0.455 0.764 0.621 0.175 0.317 0.506 0.702 0.632 0.262 0.427
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 10 0.322 0.456 0.769 0.602 0.165 0.313 0.495 0.699 0.553 0.258 0.386
LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 25 0.247 0.319 0.864 0.604 0.152 0.306 0.493 0.682 0.597 0.243 0.408
LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 25 0.241 0.315 0.872 0.582 0.141 0.298 0.477 0.683 0.523 0.236 0.370
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 25 0.246 0.316 0.867 0.603 0.148 0.305 0.492 0.682 0.592 0.242 0.406
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 25 0.240 0.311 0.876 0.581 0.138 0.298 0.475 0.682 0.523 0.235 0.369
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 25 0.245 0.314 0.870 0.603 0.150 0.306 0.493 0.682 0.592 0.242 0.406
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 25 0.239 0.310 0.878 0.581 0.140 0.298 0.476 0.682 0.525 0.236 0.370
LoRA-MCL ($\varepsilon=0.0$) BS 1 0.496 0.707 0.211 0.649 0.174 0.323 0.518 0.702 0.648 0.255 0.428
LoRA-MCL ($\varepsilon=0.0005$) BS 1 0.450 0.678 0.292 0.654 0.195 0.333 0.530 0.710 0.679 0.272 0.448
LoRA-MCL ($\varepsilon=0.05$) BS 1 0.438 0.663 0.326 0.640 0.198 0.332 0.517 0.710 0.611 0.272 0.414
LoRA-MCL ($\varepsilon=0.0$) BS 2 0.459 0.672 0.233 0.650 0.196 0.330 0.529 0.702 0.679 0.262 0.447
LoRA-MCL ($\varepsilon=0.0005$) BS 2 0.408 0.629 0.348 0.658 0.213 0.336 0.539 0.711 0.716 0.276 0.470
LoRA-MCL ($\varepsilon=0.05$) BS 2 0.397 0.616 0.384 0.632 0.208 0.336 0.524 0.710 0.629 0.274 0.424
LoRA-MCL ($\varepsilon=0.0$) BS 5 0.453 0.657 0.254 0.656 0.211 0.331 0.532 0.702 0.697 0.262 0.456
LoRA-MCL ($\varepsilon=0.0005$) BS 5 0.403 0.615 0.372 0.661 0.229 0.339 0.545 0.712 0.744 0.278 0.484
LoRA-MCL ($\varepsilon=0.05$) BS 5 0.392 0.599 0.419 0.640 0.221 0.337 0.530 0.710 0.660 0.274 0.442
Evaluating with Sampling-based decoding. We also report results using sampling-based decoding methods in Tables 7 and 8 for Clotho and AudioCaps. Specifically, we use Top-$k$ sampling with $k=50$, Top-$p$ sampling with $p=0.95$, and Typical sampling with a threshold of $0.95$. For all sampling methods, we apply a repetition penalty of $1.1$, following [29]. In these experiments, the temperature was set to $\eta=1.0$.
We observe that both LoRA-MLE and LoRA-MCL yield significantly higher diversity compared to MAP decoding, albeit at the cost of reduced output quality. This trade-off is evident when comparing results in Tables 5 and 6. While both methods show an overall improvement in quality on the CIDEr and SPIDEr metrics across datasets, performance on other metrics such as METEOR and SPICE remains relatively comparable. In terms of diversity, there does not appear to be a clear trend indicating which method performs best, as the conclusions vary depending on the dataset. These findings suggest the need for further evaluation with different values of epsilon to better understand the quality-diversity trade-off in sampling-based decoding. Additionally, a more in-depth exploration of how the number of generated hypotheses affects sampling quality—and its implications for test-time inference scaling with LoRA-MCL [85]—is left for future work.
Table 7: Results for Clotho with 5 candidates and Sampling-based decoding.

### v2

@@ E.6.1 Experimental setup
Architecture and training. We used the Instructed version of Qwen-2-Audio (Chu et al., 2024) as the base model, which features $\sim 8.4$ billion parameters. We trained using bfloat16 precision. We used LoRA adapters applied to the $Q,K,V$ linear projections of the attention modules, and the upside and downside projections of the feedforward blocks, for all the transformer blocks, which include both the audio encoder and language model decoder. We used a rank $r$, with $r=\alpha=8$ unless otherwise stated, with dropout equal to $0.1$ (enabled during training, and disabled during inference). We trained with a batch size of $2$, with AdamW optimizer (Loshchilov and Hutter, 2017) (with $\beta_{1}=0.9$, and $\beta_{2}=0.98$), weight decay of $0.05$, using a cosine scheduler with minimum learning rate of $10^{-6}$ and maximum learning rate of $10^{-5}$, with a warmup ratio of $0.1$. Gradient clipping is used with a maximum gradient norm of $1.0$. The validation loss was computed once at the end of the epoch.
E.6.2 Ablation on the relaxation parameters in LoRA-MCL
Effect of $\varepsilon$. Let $\ell_{k}=-\log p\left(x\,|\,c,\theta_{k}\right)$ denote the NLL of hypothesis $k$ for a pair $(c,x)$. Recall that the relaxed MCL loss for such a pair is
$(1-\varepsilon)\ell_{k}^{\star}+\frac{\varepsilon}{K-1}\sum_{k=1,k\neq k^{\star}}^{K}\ell_{k}=\left(1-\frac{K\varepsilon}{K-1}\right)\ell_{k}^{\star}+\frac{\varepsilon}{K-1}\sum_{k=1}^{K}\ell_{k}\;,$
where $k^{\star}(x,c)=\mathrm{argmax}_{k}\;p(x\,|\,c;\theta_{k})$ is the winner hypothesis. We see that the first term is a force that pushes the hypothesis $k$ toward the winner hypothesis, while the second term assigns equal weight to each hypothesis, pulling them toward the barycenter of the conditional distribution (See Figure 1 in (Perera et al., 2024)). The first term vanishes when $\varepsilon=\frac{K-1}{K}$, which provides an upper bound on the value of $\varepsilon$. In practice, when choosing $\varepsilon$ for a new task, the smaller the better unless a collapse is observed; we recommend trying $\varepsilon\in[0,0.1]$. We run LoRA-MCL (relaxed) with $\varepsilon\in\{0.0005,0.05,0.1,0.3,0.5,0.8\}$ on AudioCaps (AC) and Clotho (CL) in Table 5 with $K=5$ (LoRA-MCL “relaxed” rows). We observe that outside the “small” $\varepsilon$ regime ($0.0005,0.05$), increasing $\varepsilon$ significantly degrades the diversity (e.g., mBLEU 4 from $0.410$ with $\varepsilon=0.0005$ on AudioCaps to $0.963$ with $\varepsilon=0.8$ ). For quality (as measured by Oracle SPIDEr), performance also tends to degrade outside the small $\varepsilon$ regime. This occurs because large $\varepsilon$ diminishes the benefits of MCL training, causing the model to behave increasingly like LoRA-MLE, where the heads eventually become uniformised.
Effect of the temperature scheduler. LoRA-MCL annealed is based on equations (4) and (5). The annealed method relies on the theoretical foundation of deterministic annealing (e.g., Rose et al. (2002); Perera et al. (2024)), which predicts phase transition phenomena in which the hypotheses “split” at specific temperature levels, namely critical temperatures, to explore hierarchically different modes of the conditional distribution. Given a fixed number of training steps, the temperature schedule needs to be neither too slow, otherwise the final training step may still occur above the critical temperature, nor too fast, in which cause the model does not fully benefit from annealing and collapse may still occur. We provide an ablation on the decay rate in Table 5 (see the annealed rows) for $\rho\in\{0.9,0.995,0.999,0.9999\}$. Performance remains good accross most of this range, except for $\rho=0.9999$ where we observe a degradation of diversity and performance. We attribute this to the final temperature remaining above the critical temperature.
Table 5: Ablation on the relaxation parameters in LoRA-MCL. $\varepsilon$ is the relaxation parameter, and $\rho$ in the temperature scheduler $\uptau(t)=\rho^{t}\uptau_{0}$, with $\uptau_{0}=1.0$. Evaluation on AudioCaps (AC) and Clotho (CL).
Training Decoding Beam $\mathrm{mBLEU_{4}}$ (AC) $\mathrm{mBLEU_{4}}$ (CL) $\mathrm{SPIDEr}$ (AC) $\mathrm{SPIDEr}$ (CL)
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 5 0.448 0.446 0.662 0.423
LoRA-MCL ($\varepsilon=0.0005$) BS 5 0.410 0.488 0.706 0.436
LoRA-MCL ($\varepsilon=0.05$) BS 5 0.491 0.478 0.728 0.434
LoRA-MCL ($\varepsilon=0.1$) BS 5 0.502 0.526 0.738 0.435
LoRA-MCL ($\varepsilon=0.3$) BS 5 0.624 0.643 0.700 0.421
LoRA-MCL ($\varepsilon=0.5$) BS 5 0.766 0.790 0.634 0.390
LoRA-MCL ($\varepsilon=0.8$) BS 5 0.963 0.952 0.508 0.329
LoRA-MCL ($\rho=0.9$) BS 5 0.458 0.432 0.701 0.429
LoRA-MCL ($\rho=0.995$) BS 5 0.422 0.490 0.729 0.438
LoRA-MCL ($\rho=0.999$) BS 5 0.423 0.478 0.716 0.443
LoRA-MCL ($\rho=0.9999$) BS 5 0.830 0.947 0.610 0.331
E.6.3 Comparison of LoRA-MCL with multi-head fine-tuning
In this section, we compare LoRA-MCL with multi-head MCL. Using Qwen-2-Audio for audio captioning, we duplicated the Language model head (“LMHead”, 640M params) $K$ times, froze the rest of the model, and applied MCL training (without LoRA) with both relaxed and annealed variants. As an additional variant, showing that parameter count alone is not the only issue, we also duplicated an internal module (the multimodal projector “MMProj”, $\sim$ 5.2M parameters) and trained it with the same scheme. As discussed in Section 3.3, multi-head initialization is not trivial. Copying the pretrained head yields no initialization diversity, while random reinitialization discards pretrained knowledge.
Results are displayed in Table 6. We find that $(i)$ Random initialization (Init “Random”) degrades quality for both MMProj and LMHead, and in the LMHead case, produces largely unintelligible captions with artificially high diversity. $(ii)$ Copying the parameters of the pretrained model (Init “Copy”) produces coherent outputs, but does not meaningfully benefit from MCL training due to amplified collapse risk, yielding significantly lower quality in practice. These findings show that multi-head approaches are not competitive with LoRA-MCL in a comparable training setup in the setting of Large language models.
Table 6: Comparison of LoRA-MCL with Multi-head fine-tuning. Comparison is done with $K=5$ hypotheses in AudioCaps, with the same experimental setup as in the paper. The “Init” column refer to the initialization technique for the trainable parameters. In the Multi-head versions, the trainable weights are either the Language model head (LMHead) or the Multimodal Projector (MMProj). The latter are trained with the relaxed variant with $\varepsilon=0.05$.
Training Init $K$ Decoding Beam $\mathrm{mBLEU_{4}}$ $\mathrm{SPIDEr}$
Multi-head (LMHead) Copy 5 BS 5 0.489 0.561
Multi-head (LMHead) Random 5 BS 5 0.001 0.002
Multi-head (MMProj) Copy 5 BS 5 0.530 0.394
Multi-head (MMProj) Random 5 BS 5 0.256 0.140
LoRA-MCL ($\varepsilon=0.05$) LoRA 5 BS 5 0.491 0.728
LoRA-MCL (annealed) LoRA 5 BS 5 0.423 0.716
E.6.4 Additional results
Evaluating with Sampling-based decoding. We also report results using sampling-based decoding methods in Tables 9 and 11 for Clotho and AudioCaps. Specifically, we use Top-$k$ sampling with $k=50$, Top-$p$ sampling with $p=0.95$, and Typical sampling with a threshold of $0.95$. For all sampling methods, we apply a repetition penalty of $1.1$, following (Keskar et al., 2019). In these experiments, the temperature was set to $\eta=1.0$.
We observe that both LoRA-MLE and LoRA-MCL yield significantly higher diversity than MAP decoding, albeit at the cost of reduced output quality. This trade-off is evident in Tables 8 and 10. On AudioCaps, LoRA-MCL shows a slight improvement in both quality and diversity compared to LoRA-MLE. On Clotho, it provides a small gain in diversity, while quality slightly favors LoRA-MLE. These findings highlight the need for further evaluation with different annealing schedules to better characterize the quality–diversity trade-off in sampling-based decoding. Moreover, a deeper study of how the number of generated hypotheses influences sampling quality and its implications for test-time inference scaling with LoRA-MCL (Zhao et al., 2025) is left for future work.
Further comparison with prior work. We compare against Zhu et al. (2025) a diffusion-based method using retrieval-guided Langevin dynamics (DAC-RLD), which we identify as the strongest published method for diverse audio captioning on Clotho and AudioCaps with publicly available code and checkpoints. A VAE-based method from Zhang et al. (2024) is also open-sourced (Clotho only) but underperforms DAC-RLD in both quality and diversity according to the reported scores. Adversarial training has also been explored (Mei et al., 2022b; Xu et al., 2022), though these methods are outperformed by DAC-RLD, and do not release code. We evaluate the pretrained DAC-RLD checkpoints using our setup: 5 candidate captions generated per audio (instead of 50 followed by Minimum Bayes Risk decoding (Tromble et al., 2008) in the original work), oracle sentence-level quality metrics, and mBLEU-4 for diversity. For DAC-RLD, we run both with Beam Search and Nucleus sampling, as in the original work. Results (Table 7) show that DAC-RLD achieves diversity comparable to LoRA-MCL with nucleus sampling (slightly better or worse depending on the dataset) but at the cost of substantially lower SPIDEr scores than LoRA-MCL on both Clotho and AudioCaps.
Table 7: Comparison against DAC-RLD (Zhu et al., 2025) in Diverse Audio Captioning. Evaluation on AudioCaps (AC) and Clotho (CL).
Training Decoding Beam $\mathrm{mBLEU}_{4}$ (AC) $\mathrm{mBLEU}_{4}$ (CL) SPIDEr (AC) SPIDEr (CL)
DAC-RLD Nucleus ($p=0.95)$ 1 0.157 0.150 0.435 0.244
DAC-RLD BS 5 0.239 0.215 0.505 0.287
LoRA-MCL (annealed) Nucleus ($p=0.95)$ 1 0.163 0.098 0.569 0.325
LoRA-MCL ($\varepsilon=0.05$) BS 5 0.491 0.478 0.728 0.434
LoRA-MCL ($\rho=0.999$) BS 5 0.423 0.478 0.716 0.443
Table 8: Results for Clotho with 5 hypotheses and MAP Decoding. ‘BS’, ‘DBS’, ‘SR’ and ‘TTA’ stand for beam search, diverse beam search, stochastic router, and test-time augmentation respectively. We refer to the TTA parameters as $\phi_{i}$, where the strength that increases with $i$; $\phi_{1}=(0.2,0.3)$, $\phi_{2}=(0.4,0.6)$, and $\phi_{3}=(0.6,0.9)$ as the time and frequency proportion mask with SpecAugment (See Apx. E.4).
@@ E.6.4 Additional results
Training Decoding Beam $\mathrm{Div}_{2}$ $\mathrm{mBLEU}_{4}$ $\mathrm{BLEU}_{1}$ $\mathrm{BLEU}_{4}$ $\mathrm{METEOR}$ $\mathrm{ROUGE}_{L}$ $\mathrm{sBERT}$ $\mathrm{CIDEr}_{D}$ $\mathrm{SPICE}$ $\mathrm{SPIDEr}$
LoRA-MLE ($r=8$) BS 5 0.365 0.822 0.656 0.137 0.228 0.445 0.575 0.626 0.174 0.394
LoRA-MLE ($r=40$) BS 5 0.367 0.818 0.643 0.115 0.226 0.435 0.570 0.595 0.172 0.376
LoRA-MLE ($r=8$) BS 10 0.391 0.783 0.661 0.142 0.236 0.453 0.578 0.646 0.181 0.406
LoRA-MLE ($r=40$) BS 10 0.397 0.777 0.647 0.127 0.232 0.445 0.573 0.615 0.177 0.388
LoRA-MLE ($r=8$) BS 25 0.405 0.759 0.658 0.144 0.236 0.455 0.579 0.648 0.182 0.407
LoRA-MLE ($r=40$) BS 25 0.415 0.746 0.648 0.131 0.234 0.447 0.578 0.625 0.181 0.395
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 5 0.605 0.446 0.686 0.147 0.241 0.471 0.601 0.678 0.193 0.423
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 5 0.613 0.440 0.677 0.140 0.239 0.463 0.599 0.659 0.194 0.414
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 5 0.625 0.417 0.685 0.148 0.242 0.470 0.602 0.681 0.194 0.425
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 5 0.634 0.407 0.678 0.142 0.239 0.463 0.600 0.660 0.193 0.414
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 10 0.487 0.712 0.670 0.143 0.238 0.460 0.594 0.656 0.191 0.414
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 10 0.499 0.694 0.661 0.132 0.235 0.448 0.590 0.634 0.187 0.401
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 10 0.491 0.708 0.671 0.143 0.238 0.456 0.595 0.662 0.192 0.417
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 10 0.501 0.696 0.659 0.131 0.234 0.444 0.591 0.629 0.188 0.397
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 25 0.368 0.818 0.653 0.135 0.228 0.444 0.575 0.626 0.176 0.394
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 25 0.374 0.811 0.644 0.115 0.226 0.435 0.571 0.594 0.174 0.377
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 25 0.368 0.818 0.652 0.134 0.228 0.443 0.574 0.622 0.174 0.392
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 25 0.372 0.812 0.643 0.115 0.226 0.435 0.571 0.592 0.174 0.376
LoRA-MLE ($r=8$) TTA BS ($\phi_{1}$) 1 0.443 0.699 0.652 0.114 0.225 0.440 0.581 0.608 0.174 0.383
LoRA-MLE ($r=8$) TTA BS ($\phi_{2}$) 1 0.540 0.558 0.663 0.130 0.230 0.453 0.588 0.634 0.179 0.397
LoRA-MLE ($r=8$) TTA BS ($\phi_{3}$) 1 0.642 0.404 0.653 0.118 0.225 0.441 0.581 0.596 0.174 0.376
LoRA-MLE ($r=8$) TTA BS ($\phi_{1}$) 5 0.412 0.745 0.655 0.137 0.232 0.452 0.580 0.637 0.181 0.402
LoRA-MLE ($r=8$) TTA BS ($\phi_{2}$) 5 0.516 0.593 0.681 0.156 0.243 0.470 0.591 0.685 0.194 0.430
LoRA-MLE ($r=8$) TTA BS ($\phi_{3}$) 5 0.619 0.445 0.675 0.148 0.236 0.463 0.586 0.648 0.186 0.407
LoRA-MoE BS 5 0.363 0.823 0.650 0.131 0.228 0.443 0.571 0.614 0.173 0.387
LoRA-MoE BS 10 0.395 0.783 0.659 0.137 0.236 0.452 0.579 0.643 0.181 0.405
LoRA-MoE BS 25 0.408 0.757 0.657 0.140 0.237 0.454 0.579 0.643 0.184 0.405
LoRA-MoE DBS ($\lambda=0.8$) 5 0.611 0.441 0.682 0.143 0.241 0.468 0.599 0.668 0.194 0.418
LoRA-MoE DBS ($\lambda=1.0$) 5 0.630 0.410 0.682 0.147 0.241 0.467 0.600 0.675 0.195 0.422
LoRA-MoE DBS ($\lambda=0.8$) 10 0.490 0.706 0.670 0.142 0.238 0.459 0.593 0.662 0.192 0.416
LoRA-MoE DBS ($\lambda=1.0$) 10 0.494 0.705 0.668 0.138 0.237 0.456 0.592 0.650 0.191 0.410
LoRA-MoE DBS ($\lambda=0.8$) 25 0.370 0.814 0.650 0.128 0.228 0.445 0.572 0.613 0.175 0.386
LoRA-MoE DBS ($\lambda=1.0$) 25 0.370 0.814 0.651 0.129 0.229 0.445 0.573 0.620 0.174 0.390
LoRA-MoE SR BS 1 0.308 0.878 0.604 0.090 0.203 0.406 0.554 0.515 0.153 0.330
LoRA-MoE SR BS 2 0.317 0.868 0.629 0.116 0.217 0.427 0.564 0.576 0.163 0.364
LoRA-MoE SR BS 5 0.309 0.880 0.620 0.109 0.217 0.424 0.562 0.561 0.162 0.357
LoRA-MCL ($\varepsilon=0.0005$) BS 1 0.605 0.440 0.675 0.135 0.233 0.458 0.597 0.654 0.187 0.408
LoRA-MCL ($\varepsilon=0.05$) BS 1 0.617 0.415 0.680 0.130 0.239 0.463 0.601 0.662 0.190 0.414
LoRA-MCL (annealed) BS 1 0.621 0.415 0.673 0.131 0.237 0.458 0.599 0.665 0.189 0.415
LoRA-MCL ($\varepsilon=0.0005$) BS 2 0.591 0.461 0.688 0.154 0.242 0.472 0.598 0.688 0.192 0.428
LoRA-MCL ($\varepsilon=0.05$) BS 2 0.593 0.452 0.692 0.160 0.247 0.481 0.599 0.694 0.193 0.431
LoRA-MCL (annealed) BS 2 0.595 0.456 0.687 0.158 0.245 0.472 0.599 0.698 0.196 0.435
LoRA-MCL ($\varepsilon=0.0005$) BS 5 0.581 0.488 0.687 0.156 0.244 0.476 0.599 0.700 0.196 0.436
LoRA-MCL ($\varepsilon=0.05$) BS 5 0.581 0.478 0.689 0.162 0.249 0.480 0.599 0.697 0.196 0.434
LoRA-MCL (annealed) BS 5 0.584 0.478 0.689 0.168 0.246 0.477 0.600 0.711 0.199 0.443
@@
Table 9: Results for Clotho with 5 hypotheses and Sampling-based Decoding.